OpenStudy (abmon98):

for the curve xy(x+y)=84 find dy/dx at (3,4) b) find the gradient of the curve x^3-2y^3-3xy at the point (2,1)

OpenStudy (abmon98):

\[x'*y*(y+x)+x*y'*(x+y)+x*y*((x+y))'\]

OpenStudy (aum):

xy(x+y)=84 x^2y + xy^2 = 84 x^2y' + 2xy + 2xyy' + y^2 = 0 y'(x^2 + 2xy) = -(y^2 + 2xy) y' = -(y^2 + 2xy) / (x^2 + 2xy) Put x = 3, y = 4 and evaluate y'

OpenStudy (abmon98):

@aum shouldnt we differentiate xy using product rule and then using sum rule?

OpenStudy (abmon98):

for the inside brackets.

OpenStudy (abmon98):

d/dx(x)*y(x+y)+dy/dx(y)*x*(x+y)+xy(d/dx(x)+dy/dx(y))

OpenStudy (aum):

I just used algebraic expansion of xy(x+y) which makes it a bit simpler to differentiate. xy(x+y) = x^2y + xy^2 Or you can keep xy(x+y) as is and differentiate. Both should yield the same result.

OpenStudy (abmon98):

Oh you're right thank you :) but can i ask one more question

OpenStudy (aum):

Sure.

OpenStudy (abmon98):

the curve y=e^x+4e^2x has one stationary point find the x-coordinate of this point. Determine whether the stationary point is maximum or minimum.

OpenStudy (aum):

Find y', equate it to zero and solve for x. You can either do a second derivative test or an increasing/ decreasing test to determine if it is a minimum or a maximum.

OpenStudy (aum):

Is it \(\Large y = e^x + 4e^{2x}\) ?

OpenStudy (aum):

or \(\Large y = e^x + 4e^{2}x\) ?

OpenStudy (aum):

I don't think either one of those curves has a stationary point. Was there some typo in the function?

OpenStudy (abmon98):

@aum \[y=e^x+4e ^{-2x}\] \[0=e^x+(4*e ^{-2x})\]\[e^x+(0*e ^{-2x}+4*e ^{-2x})\]\[-2*4*e ^{-2x}=-e^x\]\[-8e ^{-2x}=-e^x\]

OpenStudy (abmon98):

Should i ln both sides or log both sides

OpenStudy (aum):

This function makes sense. You left out the negative sign before 4e^(-2x)

OpenStudy (abmon98):

am sorry :/

OpenStudy (aum):

But no need to apply product rule when it is a constant. d/dx (4 * e^(-2x)) = 4 * d/dx (e^(-2x)) = -8e^(-2x) -8e^(-2x) = -e^x 8e^(-2x) = e^x 8e^(-2x) / e^x = 1 8e^(-2x - x) = 1 8e^(-3x) = 1 8 = e^(3x) 3x = ln(8) x = ln(8) / 3

OpenStudy (aum):

x = ln(8) / 3 = ln(2^3) / 3 = 3ln(2) / 3 = ln(2)

OpenStudy (abmon98):

I dont get this part 8e^(-2x) = e^x Did you divide by e^x

OpenStudy (aum):

Isn't that what you had before after taking the derivative and equating it to zero? I started from that. Look at the last line in your earlier post.

OpenStudy (abmon98):

oh now i get you divided by -e^x

OpenStudy (aum):

yes.

OpenStudy (abmon98):

thank you so much :D

OpenStudy (aum):

You are welcome.

OpenStudy (abmon98):

on the maximum and minimum part we take the second derivative as you said